OpenStudy (anonymous):

Find the maximum and minimum of 6x-8y

OpenStudy (kainui):

So what is 6x-8y? Are you sure you wrote the right thing?

OpenStudy (anonymous):

\[f(x)=6x-8y\] \[f'(x)=6-8dy/dx\] \[f''(x)=-8 <0\] You have a maximum at f(x) - actually there must have been a chance of substituting value in x in \[f''(x)\]

OpenStudy (anonymous):

The proper steps are, \[f(x) = \alpha\] the differentiate the function, then equate it with zero. So you have a value of x. Now double differentiate the original function, and substitute the value of x you got in it. If the resultant value after substitution is <0 , then its a maximum at x , else its a minimum at x.

OpenStudy (kainui):

@Talshiar Those are some pretty nice assumptions you're making. But you know what they say about assumptions...

OpenStudy (anonymous):

I think so the question posted here is incomplete, or without any conditions mentioned so i assumed it to be a function of x and differentiated it.

OpenStudy (kainui):

Yeah that's why I asked to begin with. I'm not telling you how you shouldn't waste your time though.